\int \:cos5xcos2xdx
1
Expert's answer
2020-10-05T11:42:20-0400

"\textsf{Let}\hspace{0.1cm}y = \cos(5x)\cos(2x)\\


\textsf{It is known that}\\

\cos(A - B ) = \cos(A)\cos(B) + \sin(A)\sin(B)\hspace{0.1cm}\& \\

\cos(A + B ) = \cos(A)\cos(B) - \sin(A)\sin(B) \\


\textsf{Adding both equations yields,}


\displaystyle\frac{\cos(A - B )}{2} + \frac{\cos(A + B)}{2}= \cos(A)\cos(B) \\


\therefore y = \frac{\cos(7x )}{2} + \frac{\cos(3x)}{2}\\

\begin{aligned}
\int y\hspace{0.1cm}\mathrm{d}x
& = \int \frac{\cos(7x )}{2} + \frac{\cos(3x)}{2} \hspace{0.1cm}\mathrm{d}x \\&= \frac{\sin(7x )}{14} + \frac{\sin(3x)}{6} + C
\end{aligned}\\



\int \cos(5x)\cos(2x) \displaystyle\hspace{0.1cm}\mathrm{d}x = \frac{\sin(7x )}{14} + \frac{\sin(3x)}{6} + C\\
\textsf{Where}\hspace{0.1cm} C\hspace{0.1cm}\textsf{is an arbitrary constant}"